the model M = -0.711+206.6 is used to determine the maximum heart rate of M of and individual who is age A
solve for A
according to this model what is the age of and individual whose maximum heart rate is 155
simpliy all answers

Answer by nerdybill(7384)

the model M = -0.711+206.6 is used to determine the maximum heart rate of M of and individual who is age A
solve for A
according to this model what is the age of and individual whose maximum heart rate is 155
.
M = -0.711A + 206.6
Where
M is "max heartrate"
A is age
.
M - 206.6 = -0.711A
(M - 206.6)/(-0.711) = A
(206.6-M)/(0.711) = A
.
Plug in 155 for M:
(206.6-155)/(0.711) = A
(51.6)/(0.711) = A
72.6 years old = A